Errors in “The gcd-sum function”

April 16, 2007

Corrections to some errors in [1]: Page 7 line 8: the second g should be Id Page 9 line 3: 5.1 should be 4.1 Page 12 line 2: 5.4 should be 4.3 page 12 line 3: 5.3 should be 4.2 Page 13 line 13 and line -1: e, d should be e.d, Page 14 line -1: 5.4 should be 4.4.

References

[1] K. A. Broughan. The gcd-sum function, J. Integer Sequences, 4 (2001), Article 01.2.2.